# Verifying Well-Posedness of Linear PDEs using Convex Optimization

### 01 Apr 2026 v1

@@ A-C Wave Equation
Consider the wave equation with mixed boundary conditions,
@@
For this equation, solutions are given by
$u(t,s)$ $:=\sum_{n=0}^{\infty}\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}(s)$
@@
$a_{n}$ $:=2\left\langle\boldsymbol{\phi}_{n},u(0,\cdot)\right\rangle_{L_{2}}=2\int_{0}^{1}\sin(\mu_{n}s)u(0,s)\,ds,$
$b_{n}$ $:=2\left\langle\boldsymbol{\phi}_{n},u_{t}(0,\cdot)\right\rangle_{L_{2}}=2\int_{0}^{1}\sin(\mu_{n}s)u_{t}(0,s)\,ds.$
It follows that
@@
$u_{s}(t,s)$ $:=\sum_{n=0}^{\infty}\mu_{n}\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}^{\prime}(s),$
where $\boldsymbol{\phi}_{n}^{\prime}(s):=\cos(\mu_{n}(s))$.
Now, to formulate the wave equation as an abstract Cauchy problem as in (3), we may introduce the state $\mathbf{u}(t)=\begin{bmatrix}u_{t}(t)\\ u_{s}(t)\end{bmatrix}$, defining
@@
$(e^{At}\mathbf{x})(s):=\sum_{n=0}^{\infty}\mu_{n}\begin{bmatrix}\bigl(-a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}(s)\\ \bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}^{\prime}(s)\end{bmatrix}$.
@@
$\left\lVert{e^{At}\mathbf{x}}\right\rVert_{L_{2}}^{2}$ $=\sum_{n=0}^{\infty}\frac{\mu_{n}^{2}}{2}\Bigl[\bigl(-a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)^{2}$
$\qquad+\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)^{2}\Bigr]$
@@
Hence, $e^{At}$ is a contraction semigroup.
@@
whence $\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}^{2}=1$ and $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}|_{t=1}=\frac{1}{2}\mu_{m}^{2}$. It follows that, for any $M\geq 1$ and $\omega\in\mathbb{R}$, we can find sufficiently large $m\in\mathbb{N}$ such that $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}|_{t=1}=\frac{1}{2}\mu_{m}^{2}>Me^{\omega}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}^{2}$. Since the $C_{0}$-semigroup generated by $A$ must satisfy $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{H}\leq Me^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{H}$ for some $\omega\in\mathbb{R}$ and $M\geq 1$, it follows that the considered $A:D\to H$ cannot be the infinitesimal generator of a $C_{0}$-semigroup, wherefore this formulation of the wave equation cannot be well-posed.

### 02 Apr 2026 v2

@@ A-C Wave Equation
Consider the following wave equation,
@@
Solutions to this PDE are given by
$u(t,s)$ $:=\sum_{n=0}^{\infty}\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}(s),$
@@
$a_{n}:=2\left\langle\boldsymbol{\phi}_{n},u(0,\cdot)\right\rangle_{L_{2}}\qquad b_{n}:=2\left\langle\boldsymbol{\phi}_{n},u_{t}(0,\cdot)\right\rangle_{L_{2}}.$
Letting $\boldsymbol{\phi}_{n}^{\prime}(s):=\cos(\mu_{n}(s))$, it follows that
@@
$u_{s}(t,s)$ $:=\sum_{n=0}^{\infty}\mu_{n}\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}^{\prime}(s).$
Now, to formulate the wave equation as in (3), we may introduce the state $\mathbf{u}(t)=\begin{bmatrix}u_{t}(t)\\ u_{s}(t)\end{bmatrix}$, defining
@@
$(e^{tA}\mathbf{x})(s):=\sum_{n=0}^{\infty}\mu_{n}\begin{bmatrix}\bigl(-a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}(s)\\ \bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}^{\prime}(s)\end{bmatrix}$,
@@
$\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}$ $=\sum_{n=0}^{\infty}\frac{\mu_{n}^{2}}{2}\Bigl[\bigl(-a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)^{2}$
$\hskip 56.9055pt+\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)^{2}\Bigr]$
@@
Hence, $e^{tA}$ is a contraction semigroup.
@@
whence $\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}^{2}=1$ and $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}|_{t=1}=\frac{1}{2}\mu_{m}^{2}$. It follows that, for any $M\geq 1$ and $\omega\in\mathbb{R}$, we can find sufficiently large $m\in\mathbb{N}$ such that $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}|_{t=1}=\frac{1}{2}\mu_{m}^{2}>Me^{\omega}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}^{2}$. Since the $C_{0}$-semigroup generated by $A$ must satisfy $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq Me^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}$ for some $\omega\in\mathbb{R}$ and $M\geq 1$, this implies that the considered $A:D\to L_{2}^{2}$ cannot be the infinitesimal generator of a $C_{0}$-semigroup, wherefore this formulation of the wave equation cannot be well-posed.